I need help finding the risk free rate and the market risk premium for microsofts corporation??? I want to use these number for calculation of dividend discount model
Answer #1

Please provide the Numbers.

Apart from it :-

The risk-free rate of return - "Rf" Is a rate of return which involves 0 ZERO risks. In practice, however, the risk-free rate does not exist because even the safest investments carry a very small amount of risk.

the interest rate on a three-month U.S. Treasury bill is often used as the risk-free rate for U.S.-based investors

Todays Rf = 2.38%

Risk Premium = Expected market return - Risk free rate of return.

Risk Free already discussed.

Expected market return =

Common uses of the required rate of return include:

  • Calculating the present value of dividend income for the purpose of evaluating stock prices ( Dividend Discount Model)
  • Calculating the present value of free cash flow to equity
  • Calculating the present value of operating free cash flow

Dividend discount model - Gordon Growth Formula

K = ( D/S ) + G

  • k = required rate of return
  • D = dividend payment (expected to be paid next year)
  • S = current stock value (if using the cost of newly issued common stock you will need to minus the flotation costs)
  • g = growth rate of the dividend

Current Stock Value = $120.95

Flotation cost = Lets say = $3

S = 120.95 - 3 = $117.95

g= We need to identify by doing the analysis of past trends of dividend.

d= (Current year dividend paid * 1+g)

K = Balancing Fig.